# Not all Chess960 positions are equally complex

### v1

@@ Abstract
We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions. Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the advantage conferred by moving first is a robust structural feature of the game. To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies. This measure decomposes into contributions from White and Black, $S_{W}$ and $S_{B}$, yielding a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$. Across the ensemble, $S_{\mathrm{tot}}$ varies by a factor of three, while $A$ spans from 2.5-2.5 to +1.8+1.8 bits, showing that some openings burden White and others Black. The mean $\langle A\rangle=-0.25$ bits indicates a slight tendency for White to face harder opening decisions. Standard chess (position #518, RNBQKBNR) exhibits above-average asymmetry (91st percentile) but typical overall complexity (47th percentile). The most complex opening is #226 (BNRQKBNR), whereas #198 (QNBRKBNR)is the most balanced, with both evaluation and asymmetry near zero. These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of the back-rank pieces can significantly alter strategic depth and competitive fairness. Remarkably, the classical starting position—despite centuries of cultural selection—lies far from the most balanced configuration.
@@ II.2.1 Asymmetry in Decision Complexity
II.2.2 Evolution of asymmetry during the game
@@ Acknowledgments
MB thanks the GothamChess and Blitzstream channels for providing valuable food for thought about chess.

### v2

@@ Abstract
We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions. Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the initiative is a robust structural feature. To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies. This measure decomposes into contributions from White and Black, $S_{W}$ and $S_{B}$, yielding a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$. Across the ensemble, $S_{\mathrm{tot}}$ varies by a factor of three, while $A$ spans from 2.5-2.5 to +1.8+1.8 bits, showing that some openings burden White and others Black. The mean $\langle A\rangle=-0.25$ bits indicates a slight tendency for White to face harder opening decisions. Standard chess (position #518, RNBQKBNR) exhibits above-average asymmetry (91st percentile) but typical overall complexity (47th percentile). The most complex opening is #226 (BNRQKBNR), whereas #198 (QNBRKBNR)is the most balanced, with both evaluation and asymmetry near zero. These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of the back-rank pieces can significantly alter strategic depth and competitive fairness. Remarkably, the classical starting position—despite centuries of cultural selection—lies far from the most balanced configuration.
@@ II.2.1 Asymmetry in Decision Complexity
II.2.2 Complexity and time to make a move
If the quantities $\Delta$ and $S(10)$ are meaningful indicators of the intrinsic complexity of a position, one naturally expects them to correlate with the time players spend deciding on a move. Intuitively, positions with a small $\Delta$ (i.e., many nearly equivalent candidate moves) or a large entropy $S(10)$ (i.e., a broad distribution of plausible continuations) should require more cognitive effort, and therefore more time, to resolve [ref referee].
To test this hypothesis, we analyze a freely available database of online games played on Lichess with move-level time stamps [28]. For each position, we compute $\Delta$ and $S(10)$ using Stockfish at depth 15, and measure the mean time $\tau$ required by players to make a move. We then study the dependence of $\tau$ on $\Delta$ and $S(10)$ (Fig. 5). The dataset consists of $1000$ games, corresponding to a total of 70 15070\,150 moves.
Figure 5: (a) Mean thinking time $\tau$ as a function of $\Delta$. (b) Mean thinking time $\tau$ as a function of the entropy $S(10)$. Data are computed from $1000$ games (total of 70 15070\,150 moves) from [28], using Stockfish at depth 15.
Since players of different strengths perceive and handle complexity differently, we account for Elo-dependent effects by adapting the reference threshold $\Delta_{0}$ as a function of the Elo rating. Specifically, we use
$\Delta_{0}=\begin{cases}100,&\text{Elo}<1400,\\[4.0pt] 100-90\,\frac{\text{Elo}-1400}{1000},&1400\leq\text{Elo}\leq 2400,\\[4.0pt] 10,&\text{Elo}>2400.\end{cases}$ (6)
This choice reflects the increasing ability of stronger players to discriminate between close alternatives and to process complex positions more efficiently.
As expected, we observe a clear decrease of the average thinking time $\tau$ with increasing $\Delta$, and a corresponding increase of $\tau$ with $S(10)$. Deviations from these trends are visible for very small $\Delta$ or very large $S(10)$. These fluctuations can be attributed to additional factors not captured by our complexity measures, such as opening preparation: opening positions often exhibit high complexity and small $\Delta$, yet moves are played rapidly due to memorization.
Despite these effects, the overall trends provide strong empirical support for $\Delta$ and $S(10)$ as relevant proxies for decision-making complexity in chess. They capture, in a quantitative manner, the intuitive notion that positions with many competing continuations or poorly differentiated evaluations are cognitively more demanding and require longer deliberation times.
II.2.3 Evolution of asymmetry during the game
@@ Acknowledgments
I thank the GothamChess and Blitzstream channels for providing valuable food for thought about chess. I also thank Giordano de Marzo for his suggestion about thinking time and complexity.